OpenStudy (anonymous):

I'LL BECOME A FAN OF ANYONE WHO CAN HELP WITH THESE NEXT FEW PROBLEMS: If the sin 45° is 0.707, then the cos _____° is 0.707.

OpenStudy (anonymous):

The sail of a boat is in the shape of a right triangle, as shown below. The sail of a boat is a right triangle with an acute base angle equal to 50 degrees and the height of the sail is 6 meters. Which expression shows the length, in meters, of the side adjacent to the 50 degree angle?

OpenStudy (anonymous):

6 sin 50° 6 cot 50° 6 by sec 50 degrees 6 by cot 50 degrees

OpenStudy (johnweldon1993):

Well sin and cos are cofunctions sin(x) = cos(90 - x) cos(x) = sin(90 - x) So if we have sin(45) that is equivalent to sin(45) = cos(90 - 45) sin(45) = cos(45)

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

We have an angle...the side opposite that angle...and we want the side adjacent to that angle... what trig function uses Opposite and Adjacent?

OpenStudy (anonymous):

Sin, Cosine and tangent? @johnweldon1993

OpenStudy (johnweldon1993):

Well tangent would be the function we would use here...since it uses opposite/adjacent... but now looking at your choices...there is not tangent option..hmm

OpenStudy (anonymous):

Sorry I'm just really confused. lol

OpenStudy (johnweldon1993):

On what part?

OpenStudy (anonymous):

Everything. Im in a hurry but i'm not understanding how the blank is going to be filled in with the cos is 0.707

OpenStudy (johnweldon1993):

it would be 45 like I put above

OpenStudy (anonymous):

oh okay.. lol sorry this just confuses me. what about the second one?

OpenStudy (anonymous):

The sail of a boat is in the shape of a right triangle, as shown below. The sail of a boat is a right triangle with an acute base angle equal to 50 degrees and the height of the sail is 6 meters. Which expression shows the length, in meters, of the side adjacent to the 50 degree angle?